OpenStudy (anonymous):

A colony of Fruit Flies triples every 30 minutes. If there are 19 683 Fruit Flies now, how many were present 2 hours ago?

OpenStudy (anonymous):

sos

Miracrown (miracrown):

What do you think we need to do here?

OpenStudy (anonymous):

This question requires some sort of time travel. We have to go back in time to solve.

OpenStudy (anonymous):

i just need the answer pls

Miracrown (miracrown):

We're not some answer machines. Our purpose is to guide a user, so that they arrive to the answer themselves.

OpenStudy (anonymous):

How many times has it tripled since 2 hours ago?

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

\[19683 \left(\frac{1}{3}\right)^4=243 \]